# Arithmetic Bohr radius and Local Banach space theory

### v1

@@ 2. Application
We now turn to part $(2)$. Suppose first that $2<s\leq\infty$ and $2\leq t\leq\infty$. In this case, $\ell_{s,t}$ is 22-convex (see [23, p.189]), and the conclusion follows directly from Theorems 1.5(2)(2), 1.7, and 1.8. It remains to consider the case $2<s\leq\infty$, $1\leq t\leq 2$. Here we apply Theorems 1.4, 1.7, and 1.8, together with the estimates
@@
(see [23, p.192]). The condition imposed on $\psi$ in part $(1)$ guarantees that $\ell_{\psi}\subseteq\ell_{2}$. On the other hand, the condition in part $(2)$ ensures that $\ell_{\psi}$ is 22-convex (see [23, p.189]). Taking these properties into account, the desired conclusion follows directly from Theorems 1.5 and 1.8, together with the preceding facts. This completes the proof. ∎
@@ 3. Preliminaries
for all $k\in\mathbb{N}$, all scalars $\mu_{j}\in\mathbb{C}$, and all $\varepsilon_{j}\in\mathbb{C}$ with $|\varepsilon_{j}|\leq 1$. The smallest such constant is denoted by $\chi({w_{k}})$ and is called the unconditional basis constant of $w_{k}$}. We say that $w_{k}$} is a $1$-unconditional basis if $\chi({w_{k}})=1$. The unconditional basis constant of the Banach space $W$ is defined by $\chi(W):=\inf\chi(\{w_{k}\})\in[1,\infty]$, the infimum taken over all unconditional basis $\{w_{k}\}$ of $W$. A Banach space $Z$ satisfying $\ell_{1}\subset Z\subset c_{0}$ (with normal inclusions) is called a Banach sequence space if the canonical unit vectors $e_{k}$} form a $1$-unconditional basis of $Z$. A Banach lattice $Z$ is said to be 22-convex if there exists a constant $C>0$ such that
@@
$P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}$ (3.1)
@@ 4. Auxiliary Lemmas
$\sup_{z\in r\Omega}\,\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})|z^{\alpha}|^{p}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,X}.$ (4.1)
@@
$\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})r^{p\alpha}\leq\lambda^{p}\left\lVert f\right\rVert^{p}_{\Omega_{2},X}.$ (4.2)
@@
[32, Theorem 14.5] Given any infinite dimensional complex Banach space $X$, there exist $x_{1},\ldots,x_{n}\in X$ for each $n\in\mathbb{N}$ such that $\left\lVert z\right\rVert_{\infty}/2\leq\left\lVert\sum_{j=1}^{n}x_{j}z_{j}\right\rVert\leq\left\lVert z\right\rVert_{Cot(X)}$ for every choice of $z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}$. Clearly, setting $z=e_{j}$, gives $\left\lVert x_{j}\right\rVert\geq 1/2$, where $e_{j}$ is the $j$-th canonical basis vector of $\mathbb{C}^{n}$.
@@ 5. Proof of the main results
holds for all $\epsilon>0$, and so (1.4) immediately follows. On the other hand, when $\left\lVert U\right\rVert<\lambda$, a lower bound of $R_{\lambda}(B_{Z},p,U)$ was established in [36, Theorem 1.1]. Consequently, the second part of the theorem follows from (1.4) together with[36, Theorem 1.1]. ∎
Proof of Theorem 1.2. We shall first prove the left-hand inequality of $(1)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
@@
$\leq\left\lVert U\right\rVert^{p}\left\lVert a_{0}\right\rVert^{p}+\sum_{m=1}^{\infty}R^{pm}\lambda^{p}\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}_{\Omega,X}$ (5.1)
@@
provided $R\leq\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{2\lambda^{p}-\left\lVert U\right\rVert^{p}}\right)^{\frac{1}{p}}$. This yields
@@
and hence, the desired bound follows. The uper bound of the right-hand inequality of $(1)$ directly follows from the fact $\cup^{\infty}_{m=1}\mathcal{PH}(^{m}\Omega,X)\subseteq\mathcal{PH}(\Omega,X)$. We now want to prove the left-hand inequality of $(2)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
@@
$\leq\left\lVert U\right\rVert^{p}\left\lVert a_{0}\right\rVert^{p}+\sum_{m=1}^{\infty}R^{pm}\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}_{\Omega,X}$ (5.2)
@@
provided $R\leq\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{\lambda^{p}-\left\lVert U\right\rVert^{p}+1}\right)^{\frac{1}{p}}$. This yields
@@
and hence, the desired bound follows. The uper bound of the right-hand inequality of $(2)$ directly follows from the fact $\cup^{\infty}_{m=1}\mathcal{PH}(^{m}\Omega,X)\subseteq\mathcal{PH}(\Omega,X)$. ∎
Proof of Theorem 1.3. We first want to prove the inequality $(1)$. By going the similar line of arguments as in the proof of Theorem 1.2, from (5), we obtain
@@
$\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}\leq 8^{p}\,\left\lVert\left\lVert f\right\rVert_{\Omega,X}\,I-{\operatorname{Re}\,}(a_{0})\right\rVert^{p}.$ (5.3)
@@
provided $R\leq\frac{1}{(1+8^{p})^{1/p}}$. This yields
@@
and hence, the desired bound follows. By going the similar line of arguments as in the proof of part $(2)$ of Theorem 1.2, using (5) and (5.3), we obtain
@@
provided $R\leq\frac{\lambda}{(\lambda^{p}+8^{p})^{1/p}}$. This yields
@@
and hence, the desired bound follows. This completes the proof. ∎
Proof of Theorem 1.4. Let $X=\mathcal{B}(\mathcal{H})$. We shall first prove the lower bound of $AP_{\lambda}(B_{Z},p,U)$. The following lower bound of $R_{\lambda}(B_{Z},p,U)$ was established in [36, Theorem 1.2] in the case where $X$ is finite dimensional:
@@
$2\,\sum_{|\alpha|=m}\left(\frac{m!}{\alpha!}\right)^{p}\,r^{p\alpha}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{B_{\ell^{n}_{q}},X}.$ (5.4)
@@
$=2^{p}\,\lambda^{p}\,\left(n^{1-\frac{1}{p}}\right)^{pm}\,\left\lVert\sum_{|\alpha|=m}\varepsilon_{\alpha}\,\frac{m!}{\alpha!}\,z^{\alpha}\right\rVert^{p}_{B_{\ell^{n}_{q}},\mathbb{C}}.$ (5.5)
We now recall the following probabilistic estimate due to Boas [15, Theorem 4]:
@@
$\leq C_{1}\,\lambda^{\frac{1}{m}}\,\left(m\,\log\,m(m!)^{2\left(1-\frac{1}{\min\{2,q\}}\right)}\,n\right)^{\frac{1}{2m}}\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}$ (5.6)
@@
$C_{1}\lambda^{\frac{1}{m}}\,(m\,\log\,m\,n)^{\frac{1}{2m}}\,\left(\frac{m}{e}\right)^{t}(2\pi m)^{\frac{t}{2m}}\left(1+O\left(\frac{1}{m}\right)\right)^{\frac{1}{2m}}\,\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}.$ (5.7)
@@
$AP_{\lambda}(B_{\ell^{n}_{q}},p,X)\leq C\,\lambda^{\frac{1}{\log\,n}}\,\,(\log\,n)^{1-\frac{1}{\min\{2,q\}}}\,\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}.$ (5.8)
@@
Proof of Theorem 1.5. We begin by recalling two standard facts concerning finite-dimensional Banach spaces with unconditional bases. Let $W=(\mathbb{C}^{n},||.||)$ be a Banach space whose canonical unit vectors $e_{k}$ form a normalized $1$-unconditional basis. Then the operator norm of the identity embedding from $W$ into $\ell^{n}_{2}$ is given by $\left\lVert Id:W\rightarrow\ell_{2}^{n}\right\rVert=\sup_{z\in B_{W}}(\sum_{m=1}^{n}|z_{m}|^{2})^{1/2}$. On the other hand, the identity embedding from $Y$ into $\ell^{n}_{1}$ satisfies $\left\lVert Id:W\rightarrow\ell_{1}^{n}\right\rVert=\sup_{z\in B_{W}}\sum_{m=1}^{n}|z_{m}|=\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}$ for all $z\in W$. We are now in a position to establish the assertions of the theorem.
@@
We shall now prove part $(2)$. By a result of [37, Theorem 1.d.5], we may assume without loss of generality that the constant $M^{(2)}(X)=1$. This implies that $\left\lVert Id:\ell_{2}^{n}\rightarrow Z_{n}\right\rVert=1$. We now recall the following important fact from [23, (5.3), p. 191]. If $W=(\mathbb{C}^{n},||.||)$ be a symmetric Banach space such that $M^{(2)}(W)=1$, then
$\left\lVert Id:W\rightarrow\ell^{n}_{2}\right\rVert=\frac{\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}}{\sqrt{n}}.$ (5.9)
The conclusion then follows by combining Theorem 1.4, the above facts for identity embeddings, and the inequality (5.9).This completes the proof. ∎
Proof of Theorem 1.6. Let $X=\mathcal{B}(\mathcal{H})$. The lower bound of $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)$ follows from Theorem 1.1 by observing the fact $\left\lVert Id:B_{\ell^{n}_{q}}\rightarrow B_{\ell^{n}_{1}}\right\rVert=n^{1-1/q}$. Now, we want to find the upper bound of $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
$\sum_{k=1}^{n}(\left\lVert c_{k}\right\rVert^{p}+\left\lVert d_{k}\right\rVert^{p})r^{p}\leq\left\lVert g\right\rVert^{p}_{B_{Z},\mathcal{B}(\mathcal{H})}$ (5.10)
@@
$\frac{1}{n}\,\sum_{k=1}^{n}r_{k}\leq 2^{1-\frac{1}{p}}n^{-\frac{1}{p}}\sup_{z\in B_{\ell^{n}_{q}}}\left\lVert z\right\rVert_{Cot(X)}.$ (5.11)
@@
$\left(\sum_{|\alpha|=m}\left\lVert U(c_{\alpha})\right\rVert^{\beta}_{Y}\right)^{\frac{1}{\beta}}\leq C^{m}\,\left\lVert Q\right\rVert_{\mathbb{D}^{n},X}$ (5.12)
@@
$R_{\lambda}(B_{Z},p,U)\geq\max\left\{R_{\frac{\lambda}{\left\lVert U\right\rVert}}(B_{Z},p,X),\,\,R_{\frac{\lambda}{\left\lVert U\right\rVert}}(B_{Z},p,Y)\right\}.$ (5.13)
In view of [36, Theorem 1.5] and (5.13), for $1\leq p<\infty$ and $1\leq q\leq\infty$, we deduce that
$R_{\lambda}(B_{Z},p,U)\geq\frac{\Phi_{1}(n,\lambda,p,q)}{S(B_{Z},B_{\ell^{n}_{q}})S(B_{\ell^{n}_{q}},B_{Z})},$ (5.14)
@@
We shall now prove the part $(2)$. Now let $X=\mathcal{B}(\mathcal{H})$. Let $P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\beta|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\in\mathcal{PH}(^{m}Z,X)$. Consider the holomorphic polynomial $H(z):=\sum_{|\alpha|=m}\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)z^{\alpha}$ on $B_{Z}$. In view of Hölder’s inequality and Theorem 1.9, we obtain
@@
$\leq 2(c^{m}_{1}+c^{m}_{2})\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{(s-1)mt-s+t}{mst}}\left\lVert P\right\rVert_{B_{Z},X}$ (5.15)
@@
$\left(\sum_{|\alpha|=m}\left\lVert U\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)z^{\alpha}\right\rVert^{p}\right)^{1/p}\leq 2(c^{m}_{1}+c^{m}_{2})\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{(s-1)mt-s+t}{mst}}\left\lVert P\right\rVert_{B_{Z},X}.$ (5.16)
@@
$\sum_{|\alpha|=m}\left(\left\lVert U(a_{\alpha})\right\rVert^{p}+\left\lVert U(b_{\alpha})\right\rVert^{p}\right)|z^{\alpha}|^{p}\leq C^{pm}\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{((s-1)mt-s+t)p}{mst}}\left\lVert P\right\rVert^{p}_{B_{Z},X}$ (5.17)
for some constant $C>0$. Using the estimate $\sum_{|\alpha|=m}1=\binom{n+m-1}{m}\leq e^{m}\left(1+n/m\right)^{m}$, from (5.17), we deduce that
$R^{m}(B_{Z},p,U)\geq\tilde{D}\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}$ (5.18)
@@
$R^{m}_{\lambda}(B_{Z},p,U)\geq\lambda^{\frac{1}{m}}\,\tilde{D}\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}.$ (5.19)
We now observe the elementary fact if $a,b>0$ and $0<p<1$, then $(a+b)^{p}\leq 2^{p}\,\max\{a^{p},b^{p}\}$. Clearly, $0<\frac{(s-1)mt-s+t}{mst}<1$. Thus, we have

### v2

@@ 2. Application
We now turn to part $(2)$. Suppose first that $2<s\leq\infty$ and $2\leq t\leq\infty$. In this case, $\ell_{s,t}$ is $2$-convex (see [23, p.189]), and the conclusion follows directly from Theorems 1.5(2)(2), 1.7, and 1.8. It remains to consider the case $2<s\leq\infty$, $1\leq t\leq 2$. Here we apply Theorems 1.4, 1.7, and 1.8, together with the estimates
@@
(see [23, p.192]). The condition imposed on $\psi$ in part $(1)$ guarantees that $\ell_{\psi}\subseteq\ell_{2}$. On the other hand, the condition in part $(2)$ ensures that $\ell_{\psi}$ is $2$-convex (see [23, p.189]). Taking these properties into account, the desired conclusion follows directly from Theorems 1.5 and 1.8, together with the preceding facts. This completes the proof. ∎
@@ 3. Preliminaries
for all $k\in\mathbb{N}$, all scalars $\mu_{j}\in\mathbb{C}$, and all $\varepsilon_{j}\in\mathbb{C}$ with $|\varepsilon_{j}|\leq 1$. The smallest such constant is denoted by $\chi({w_{k}})$ and is called the unconditional basis constant of $w_{k}$}. We say that $w_{k}$} is a $1$-unconditional basis if $\chi({w_{k}})=1$. The unconditional basis constant of the Banach space $W$ is defined by $\chi(W):=\inf\chi(\{w_{k}\})\in[1,\infty]$, the infimum taken over all unconditional basis $\{w_{k}\}$ of $W$. A Banach space $Z$ satisfying $\ell_{1}\subset Z\subset c_{0}$ (with normal inclusions) is called a Banach sequence space if the canonical unit vectors $e_{k}$} form a $1$-unconditional basis of $Z$. A Banach lattice $Z$ is said to be $2$-convex if there exists a constant $C>0$ such that
@@
(3.1) $P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}$
@@ 4. Auxiliary Lemmas
(4.1) $\sup_{z\in r\Omega}\,\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})|z^{\alpha}|^{p}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{\Omega,X}.$
@@
(4.2) $\sum_{m=0}^{\infty}\sum_{|\alpha|=m}(\left\lVert U(a_{\alpha})\right\rVert^{p}_{Y}+\left\lVert U(b_{\alpha})\right\rVert^{p}_{Y})r^{p\alpha}\leq\lambda^{p}\left\lVert f\right\rVert^{p}_{\Omega_{2},X}.$
@@
[33, Theorem 14.5] Given any infinite dimensional complex Banach space $X$, there exist $x_{1},\ldots,x_{n}\in X$ for each $n\in\mathbb{N}$ such that $\left\lVert z\right\rVert_{\infty}/2\leq\left\lVert\sum_{j=1}^{n}x_{j}z_{j}\right\rVert\leq\left\lVert z\right\rVert_{Cot(X)}$ for every choice of $z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}$. Clearly, setting $z=e_{j}$, gives $\left\lVert x_{j}\right\rVert\geq 1/2$, where $e_{j}$ is the $j$-th canonical basis vector of $\mathbb{C}^{n}$.
@@ 5. Proof of the main results
holds for all $\epsilon>0$, and so (1.4) immediately follows. On the other hand, when $\left\lVert U\right\rVert<\lambda$, the following lower bound of $R_{\lambda}(B_{Z},p,U)$ was established in [37, Theorem 1.1]:
$R_{\lambda}(B_{Z},p,U)\geq D.\frac{1}{\sup_{z\in B_{Z}}\left\lVert z\right\rVert_{p}},$
where $D$ as in the statement of Theorem 1.1. Consequently, the second part of the theorem follows from (1.4) together with the preceding bound. This completes the proof. ∎
Proof of Theorem 1.2. We shall first prove the left-hand inequality in $(1)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
@@
(5.1) $\leq\left\lVert U\right\rVert^{p}\left\lVert a_{0}\right\rVert^{p}+\sum_{m=1}^{\infty}R^{pm}\lambda^{p}\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}_{\Omega,X}$
@@
provided
$R\leq\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{2\lambda^{p}-\left\lVert U\right\rVert^{p}}\right)^{\frac{1}{p}}.$
This yields
@@
and hence, the desired bound follows. The uper bound of the right-hand inequality of $(1)$ directly follows from the fact $\cup^{\infty}_{m=1}\mathcal{PH}(^{m}\Omega,X)\subseteq\mathcal{PH}(\Omega,X)$.
We next prove the left-hand inequality in $(2)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
@@
(5.2) $\leq\left\lVert U\right\rVert^{p}\left\lVert a_{0}\right\rVert^{p}+\sum_{m=1}^{\infty}R^{pm}\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}_{\Omega,X}$
@@
provided
$R\leq\left(\frac{\lambda^{p}-\left\lVert U\right\rVert^{p}}{\lambda^{p}-\left\lVert U\right\rVert^{p}+1}\right)^{\frac{1}{p}}.$
This yields
@@
and hence, the desired bound follows. The uper bound of the right-hand inequality in $(2)$ directly follows from the fact $\cup^{\infty}_{m=1}\mathcal{PH}(^{m}\Omega,X)\subseteq\mathcal{PH}(\Omega,X)$. ∎
Proof of Theorem 1.3. We first prove the inequality $(1)$. By going the similar line of arguments as in the proof of Theorem 1.2, from (5), we obtain
@@
(5.3) $\left\lVert\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\alpha|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\right\rVert^{p}\leq 8^{p}\,\left\lVert\left\lVert f\right\rVert_{\Omega,X}\,I-{\operatorname{Re}\,}(a_{0})\right\rVert^{p}.$
@@
provided
$R\leq\frac{1}{(1+8^{p})^{1/p}}.$
This yields
@@
and hence, the desired bound follows.
By going the similar line of arguments as in the proof of part $(2)$ of Theorem 1.2, using (5) and (5.3), we obtain
@@
provided
$R\leq\frac{\lambda}{(\lambda^{p}+8^{p})^{1/p}}.$
This yields
@@
and hence, the desired bound follows. The proof is now complete. ∎
Proof of Theorem 1.4. Let $X=\mathcal{B}(\mathcal{H})$. We shall first prove the lower bound of $AP_{\lambda}(B_{Z},p,X)$. The following lower bound of $R_{\lambda}(B_{Z},p,X)$ was established in [37, Theorem 1.2] in the case where $X$ is finite dimensional:
@@
(5.4) $2\,\sum_{|\alpha|=m}\left(\frac{m!}{\alpha!}\right)^{p}\,r^{p\alpha}\leq\lambda^{p}\,\left\lVert f\right\rVert^{p}_{B_{\ell^{n}_{q}},X}.$
@@
(5.5) $=2^{p}\,\lambda^{p}\,\left(n^{1-\frac{1}{p}}\right)^{pm}\,\left\lVert\sum_{|\alpha|=m}\varepsilon_{\alpha}\,\frac{m!}{\alpha!}\,z^{\alpha}\right\rVert^{p}_{B_{\ell^{n}_{q}},\mathbb{C}}.$
We now recall the following probabilistic estimate due to Boas [16, Theorem 4]:
@@
(5.6) $\leq C_{1}\,\lambda^{\frac{1}{m}}\,\left(m\,\log\,m(m!)^{2\left(1-\frac{1}{\min\{2,q\}}\right)}\,n\right)^{\frac{1}{2m}}\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}$
@@
(5.7) $C_{1}\lambda^{\frac{1}{m}}\,(m\,\log\,m\,n)^{\frac{1}{2m}}\,\left(\frac{m}{e}\right)^{t}(2\pi m)^{\frac{t}{2m}}\left(1+O\left(\frac{1}{m}\right)\right)^{\frac{1}{2m}}\,\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}.$
@@
(5.8) $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)\leq C\,\lambda^{\frac{1}{\log\,n}}\,\,(\log\,n)^{1-\frac{1}{\min\{2,q\}}}\,\,\left(\frac{1}{n}\right)^{\frac{1}{p}+\frac{1}{\max\{2,q\}}-\frac{1}{2}}.$
@@
Proof of Corollary 1.4. Let $X=\mathcal{B}(\mathcal{H})$. The following lower bound of $R_{\lambda}(B_{\ell^{n}_{q}},p,X)$ was established in [37, Corollary 1.2] in the case where $X$ is finite dimensional:
$R_{\lambda}(B_{\ell^{n}_{q}},p,X)\geq\begin{cases}E^{\prime}_{3}(X)\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,\left(\frac{\log n}{n}\right)^{\left(1-\frac{1}{\min\{q,2\}}\right)}&\text{for $p=1$},\\[5.69054pt] E^{\prime}_{2}\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,n^{-\frac{1}{p}}&\text{for $p\geq q$},\\[5.69054pt] E_{4}(X)\,\left(\frac{\lambda^{p}-1}{2\lambda^{p}-1}\right)^{\frac{1}{p}}\,n^{-\frac{p-1}{p(q-1)}}\,\left(\frac{\log n}{n}\right)^{\left(1-\frac{1}{\min\{q,2\}}\right)\,\frac{q-p}{p(q-1)}}&\text{for $1<p<q$}.\end{cases}$
Here, $E^{\prime}_{3}(X)$, $E^{\prime}_{2}$, and $E_{4}(X)$ are positive constants: $E^{\prime}_{3}(X)$ depends only on $X$, $E_{4}(X)$ depends on both $X$ and $p$, while $E^{\prime}_{2}$ is independent of $X$ and depends only on $p$. Thus, by combining Theorem 1.1 with the preceding estimate of $R_{\lambda}(B_{\ell^{n}_{q}},p,X)$, the desired lower bound of $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)$ follows. The proof is now complete. ∎
Proof of Theorem 1.5. We start by recalling two classical facts from the theory of finite-dimensional Banach spaces with unconditional bases. Let $W=(\mathbb{C}^{n},||.||)$ be a Banach space whose canonical unit vectors $e_{k}$ form a normalized $1$-unconditional basis. Then the operator norm of the identity embedding from $W$ into $\ell^{n}_{2}$ is given by $\left\lVert Id:W\rightarrow\ell_{2}^{n}\right\rVert=\sup_{z\in B_{W}}(\sum_{m=1}^{n}|z_{m}|^{2})^{1/2}$. On the other hand, the identity embedding from $Y$ into $\ell^{n}_{1}$ satisfies $\left\lVert Id:W\rightarrow\ell_{1}^{n}\right\rVert=\sup_{z\in B_{W}}\sum_{m=1}^{n}|z_{m}|=\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}$ for all $z\in W$. We are now in a position to establish the assertions of the theorem.
@@
We shall now prove part $(2)$. By a result of [39, Theorem 1.d.5], we may assume without loss of generality that the constant $M^{(2)}(X)=1$. This implies that $\left\lVert Id:\ell_{2}^{n}\rightarrow Z_{n}\right\rVert=1$. We now recall the following important fact from [23, (5.3), p. 191]. If $W=(\mathbb{C}^{n},||.||)$ be a symmetric Banach space such that $M^{(2)}(W)=1$, then
(5.9) $\left\lVert Id:W\rightarrow\ell^{n}_{2}\right\rVert=\frac{\left\lVert\sum_{m=1}^{n}e^{*}_{m}\right\rVert_{W^{*}}}{\sqrt{n}}.$
The conclusion then follows by combining Theorem 1.4, the above facts for identity embeddings, and the inequality (5.9). This concludes the proof. ∎
Proof of Theorem 1.6. Let $X=\mathcal{B}(\mathcal{H})$. The lower bound of $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)$ follows from Theorem 1.1 by observing the fact
$\left\lVert Id:B_{\ell^{n}_{q}}\rightarrow B_{\ell^{n}_{1}}\right\rVert=n^{1-\frac{1}{q}}$
for any $1\leq q\leq\infty$. Now, we want to find the upper bound of $AP_{\lambda}(B_{\ell^{n}_{q}},p,X)$. Let $r=(r_{1},.\,.\,.,r_{n})\in\mathbb{R}^{n}_{\geq 0}$ be such that
(5.10) $\sum_{k=1}^{n}(\left\lVert c_{k}\right\rVert^{p}+\left\lVert d_{k}\right\rVert^{p})r^{p}\leq\left\lVert g\right\rVert^{p}_{B_{Z},\mathcal{B}(\mathcal{H})}$
@@
(5.11) $\frac{1}{n}\,\sum_{k=1}^{n}r_{k}\leq 2^{1-\frac{1}{p}}n^{-\frac{1}{p}}\sup_{z\in B_{\ell^{n}_{q}}}\left\lVert z\right\rVert_{Cot(X)}.$
@@
(5.12) $\left(\sum_{|\alpha|=m}\left\lVert U(c_{\alpha})\right\rVert^{\beta}_{Y}\right)^{\frac{1}{\beta}}\leq C^{m}\,\left\lVert Q\right\rVert_{\mathbb{D}^{n},X}$
@@
(5.13) $R_{\lambda}(B_{Z},p,U)\geq\max\left\{R_{\frac{\lambda}{\left\lVert U\right\rVert}}(B_{Z},p,X),\,\,R_{\frac{\lambda}{\left\lVert U\right\rVert}}(B_{Z},p,Y)\right\}.$
In view of [37, Theorem 1.5] and (5.13), for $1\leq p<\infty$ and $1\leq q\leq\infty$, we deduce that
(5.14) $R_{\lambda}(B_{Z},p,U)\geq\frac{\Phi_{1}(n,\lambda,p,q)}{S(B_{Z},B_{\ell^{n}_{q}})S(B_{\ell^{n}_{q}},B_{Z})},$
@@
We shall now prove the part $(2)$. Now let $X=\mathcal{B}(\mathcal{H})$. Let
$P(z)=\sum_{|\alpha|=m}a_{\alpha}\,z^{\alpha}+\sum_{|\beta|=m}b^{*}_{\alpha}\,\bar{z}^{\alpha}\in\mathcal{PH}(^{m}Z,X).$
Consider the holomorphic polynomial
$H(z):=\sum_{|\alpha|=m}\left(\frac{a_{\alpha}+b_{\alpha}}{2}\right)z^{\alpha}$
on $B_{Z}$. In view of Hölder’s inequality and Theorem 1.9, we obtain
@@
(5.15) $\leq 2(c^{m}_{1}+c^{m}_{2})\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{(s-1)mt-s+t}{mst}}\left\lVert P\right\rVert_{B_{Z},X}$
@@
(5.16) $\left(\sum_{|\alpha|=m}\left\lVert U\left(\frac{a_{\alpha}-b_{\alpha}}{2}\right)z^{\alpha}\right\rVert^{p}\right)^{1/p}\leq 2(c^{m}_{1}+c^{m}_{2})\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{(s-1)mt-s+t}{mst}}\left\lVert P\right\rVert_{B_{Z},X}.$
@@
(5.17) $\sum_{|\alpha|=m}\left(\left\lVert U(a_{\alpha})\right\rVert^{p}+\left\lVert U(b_{\alpha})\right\rVert^{p}\right)|z^{\alpha}|^{p}\leq C^{pm}\bigg(\sum_{|\alpha|=m}1\bigg)^{\frac{((s-1)mt-s+t)p}{mst}}\left\lVert P\right\rVert^{p}_{B_{Z},X}$
for some constant $C>0$. Using the estimate
$\sum_{|\alpha|=m}1=\binom{n+m-1}{m}\leq e^{m}\left(1+\frac{n}{m}\right)^{m},$
from (5.17), we deduce that
(5.18) $R^{m}(B_{Z},p,U)\geq\tilde{D}\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}$
@@
(5.19) $R^{m}_{\lambda}(B_{Z},p,U)\geq\lambda^{\frac{1}{m}}\,\tilde{D}\left(1+\frac{n}{m}\right)^{-\frac{(s-1)mt-s+t}{mst}}.$
We now observe the elementary fact if $a,b>0$ and $0<p<1$, then $(a+b)^{p}\leq 2^{p}\,\max\{a^{p},b^{p}\}$. Clearly,
$0<\frac{(s-1)mt-s+t}{mst}<1.$
Thus, we have